A lease agreement that qualifies as a finance lease calls for annual lease payments of $20,000 over a eight-year lease term (also the asset’s useful life), with the first payment at January 1, the beginning of the lease. The interest rate is 4.

a. Determine the present value of the lease upon the lease's inception.
b. Create a partial amortization through the first payment on January 1, 2017.
c. If the lessee’s fiscal year is the calendar year, what would be the pretax amounts related to the lease that the lessee would report in its income statement for the first year ended December 31 (ignore taxes)?

If you can show me the steps in a ti 84 calculator that would be Great. ****

Determine the present value of the lease upon the lease's inception. (Round your answers to nearest whole number and round percentage answer to 1 decimal place.)

PV factors based on
Table or Calculator function:
Lease Payment
n =
i =
PV of Lease

2 Create a partial amortization through the first payment on January 1, 2017. (Enter all amounts as positive values. Round your answers to nearest whole number.)

Date Lease Payment Effective Interest Decrease in balance Outstanding balance
01/01/2016
01/01/2016
01/01/2017

3 )If the lessee’s fiscal year is the calendar year, what would be the pretax amounts related to the lease that the lessee would report in its income statement for the first year ended December 31? (Round your answers to nearest whole number.)

Impact on pretax income related to the lease:
Total expenses
Answer #1
Answer a:
Present value of lease option
PV Factors based on
Table or calculate function 7.00205 Table value 7.00205
Lease payment $20,000.00
n= 8
i= 4%
PV of lease $140,041.00 =PV(4%,8,20000,,1)
or
=20000*7.00205
Answer b:
Lease Amortization Schedule
Date Lease Payment Effective Interest Decrease in balance Outstanding balance
1-Jan-16 $140,041
1-Jan-16 $20,000 $0 $20,000 $120,041
1-Jan-17 $20,000 $4,802 $15,198 $104,843
Note:- Effective interest = 120,041*4% = 4,802
Answer c:
Impact on pretax income related to lease
Particulars Amount
Interest Expense $4,802
Depreciation ($140,041/8) $17,505
Total expenses $22,307